OpenStudy (anonymous):

Two large tanks, each holding 50L of liquid, are interconnected by a pipe with liquid flowing from tank A into tank B at a rate of 5 L/min. The liquid in each tank is kept well stirred. Pure water flows into tank A at a rate of 5L/min. The solution flows out of the system from tank B at 5L/min. If, initially, tank A contains 50kg if salt and tank B contains 100kg, determine the mass of salt in each tank at time t>=0.

OpenStudy (anonymous):

Here is my complete work: let dQ/dt=the rate of change of salt in the tank at t>=0 and thus dQ/dt=rate in-rate out For Tank A, rate in=(o kg/L)(5L/min) =0 since salt is not entering the tank rate out=(QA(t) kg/50 L) x (5L/min) = QA(t)/10 kg/min dQA(t)/dt=-QA(t)/10kg/min ?????????? For Tank B, rate in=(QA(t)/50L)(5L/min) = QA(t) kg/ 10 min rate out=(QB(t)/50)(5L/min)=QB(t)/10 kg/min So, dQB(t)/dt=QA(t)/10-QB(t)/10 ????????? where QA(0)=50 and QB(0)=100